OpenStudy (anonymous):

Solve the system of Equations y = 3x + 2 x - 2y = 10 Please someone help I am a little confused

OpenStudy (anonymous):

Do you want to do substitution or elimination?

OpenStudy (anonymous):

It doesn't specify it just reads to solve the system of equations. I am sorry

OpenStudy (anonymous):

Well, there are different ways to solve it, graphing, substitution or elimination

OpenStudy (anonymous):

for this one, i'll do substitution because it's easier

OpenStudy (anonymous):

sounds good. just not to sure how to do that though but easier the better

OpenStudy (anonymous):

Since the first equation already tells you that y = 3x+2, plug in the y equation into the second equation. x -2 (3x+2)=10

OpenStudy (anonymous):

So I basically shoved the first equation into a variable for the 2nd equation

OpenStudy (anonymous):

y = 3x + 2 x - 2y = 10 x-2(3x+2) = 10 x-6x -4 = 10 -5x - 4 = 10 -5x = 14 x = (14/-5) y = 3(14/-5) + 2 y = -6.4 ---Heres the work for substitution.

OpenStudy (anonymous):

Plug the answers for each variable into the equations above, if the 2 equations presented above match, then it is correct.

OpenStudy (anonymous):

yea. that is the way to check the work :) 3( 14/-5) + 2 = ???? hopefully -6.4 and (14/-5) - 2(-6.4) = ???? hopefully 10

OpenStudy (anonymous):

I am lost

OpenStudy (anonymous):

3(14/-5)+2 = y. 3(14/-5) = -6.4, so y = -6.4. So now use the second equation and plug in those variables (14/-5)-(2 times -6.4), and if it equals 10, then it is correct

OpenStudy (anonymous):

y = 3x + 2 (Given) x - 2y = 10 (Given) x-2(3x+2) = 10 (substitued Y from first equation into second) x-6x -4 = 10 (solving for x) -5x - 4 = 10 (combined x terms) -5x = 14 (add 4 both sides) x = (14/-5) (divide by neg 5) y = 3(14/-5) + 2 (now I know what x is, substitute that back in the other equation) y = -6.4 (solved by using arithmetic)

OpenStudy (anonymous):

Yes, you are correct for the variables and the work looks good! =)

OpenStudy (anonymous):

NOTE Adjustment to this line: x-6x -4 = 10 (solving for x: used the distribution property to simply into like terms)

OpenStudy (anonymous):

Thank you so much for all of your help I understand it now Thanks again

OpenStudy (anonymous):

Np, also if you're really sick of doing substitution or elimination, you can just use geoalgebra( a free graphing system) to graph and find where the line intersects, the intersection point will represent the (x,y) values. ( not good for decimals tho)

OpenStudy (anonymous):

Besides that, np and goodluck =)

OpenStudy (anonymous):

Is the concept clear for you? (I can elaborate on what @Schen421 said if you'd like)